### v1

# Characterizing pedestrian contact interaction trajectories to understand spreading risk in human crowds

Abstract
A spreading process can be observed when particular information, substances, or diseases spread through a population over time in social and biological systems. It is widely believed that contact interactions among individual entities play an essential role in the spreading process. Although contact interactions are often influenced by geometrical conditions, little attention has been paid to understand their effects, especially on contact duration among pedestrians. To examine how the pedestrian flow setups affect contact duration distribution, we have analyzed trajectories of pedestrians in contact interactions collected from pedestrian flow experiments of uni-, bi- and multi-directional setups. Based on turning angle entropy and efficiency, we have classified the type of motion observed in the contact interactions. We have found that the majority of contact interactions in the unidirectional flow setup can be categorized as confined motion, hinting at the possibility of long-lived contact duration and hence high spreading risk. However, ballistic motion is more frequently observed in the other flow conditions, yielding frequent, brief contact interactions. Our results demonstrate that observing more confined motions is likely associated with the increase of parallel contact interactions regardless of pedestrian flow setups. This study highlights that the confined motions tend to yield longer contact duration, potentially leading to higher risk of virus exposure in human face-to-face interactions. These results have important implications for crowd management in the context of minimizing spreading risk.
This work is an extended version of Kwak et al. (2023) presented at the 2023 International Conference on Computational Science (ICCS).
keywords:
[…]
1 Introduction
Modeling contact interactions among individual entities is essential to understand spreading processes in social and biological systems, such as information diffusion in human populations [1, 2] and transmission of infectious disease in animal and human groups [3, 4]. For the spreading processes in social and biological systems, one can observe a contact interaction when two individual entities are within a close distance, so they can exchange substance and information or transmit disease from one to the other one. In previous studies, macroscopic patterns of contact interactions are often estimated based on simple random walking behaviors, including ballistic motion. For example, Rast [5] simulated continuous-space-time random walks based on ballistic motion of non-interacting random walkers. Although such random walk models have widely applied to estimate contact duration for human contact networks, little work has been done to study the influence of pedestrian flow geometrical conditions on the distribution of contact duration.
To examine how the geometrical conditions of pedestrian flow affect the contact duration distribution, we perform trajectory analysis for the experimental dataset collected from a series of experiments performed for various pedestrian flow setups. The trajectory analysis of moving organisms, including proteins in living cells, animals in nature, and humans, has been a popular research topic in various fields such as biophysics [6, 7, 8], movement ecology [9, 10, 11], and epidemiology [12, 13]. Single particle tracking (SPT) analysis, a popular trajectory analysis approach frequently applied in biophysics and its neighboring disciplines, characterizes the movement dynamics of individual entities based on observed trajectories [7, 14]. According to SPT analysis, one can identify different types of diffusion, for instance, directed diffusion in which individuals move in a clear path and confined diffusion in which individuals tend to move around the initial position. The most common method for identifying diffusion types is based on the mean-squared displacement (MSD), which reflects the deviation of an individual’s position with respect to the initial position after time lag [14, 15]. Motion types can be identified based on the diffusion exponent. MSD has been widely applied for various trajectory analysis studies in biophysics [16, 17]. For pedestrian flow trajectory analysis, Murakami et al. [18, 19] analyzed experimental data of bidirectional pedestrian flow and reported diffusive motion in individual movements perpendicular to the flow direction. They suggested that uncertainty in predicting neighbors’ future motion contributes to the appearance of diffusive motion in pedestrian flow.
Previous studies have demonstrated the usefulness of SPT analysis in examining the movement of individuals. However, SPT analysis does not explicitly consider relative motion among individuals in contact, suggesting that analyzing the relative motion can reveal patterns that might not be noticeable from the SPT analysis approach. For example, if two nearby individuals are walking in parallel with a similar speed, one might be able to see various shapes of relative motion trajectories, although the individual trajectories are nearly straight lines. For contact interaction analysis, the analysis of relative motion trajectories can be utilized to predict the length of contact duration and identify contact interaction characteristics, such as when the interacting individuals change walking direction significantly. Regarding the spreading processes, examining relative motion trajectories can be used to identify optimal geometrical conditions that can minimize contact duration when diseases are being transmitted through face-to-face (direct) interactions.
Although MSD is simple to apply, MSD has limitations. Refs. [20, 21, 22] pointed out that MSD might not be suitable for short trajectories to extract meaningful information. Additionally, due to its power-law form, the estimation of the diffusion exponent in MSD is prone to estimation errors [23, 24]. As an alternative to MSD, various approaches have been proposed, including the statistical test approach [20, 21, 25, 26] and machine learning approaches [27].
In our resent work [28], we analyzed pedestrian contact interaction trajectories based on a statistical testing approach. For the statistical test procedure, we measured a standardized value of the largest distance traveled by individuals from their starting point during the contact interaction. Ref. [28] demonstrated potential in identifying different types of contact interaction observed from experimental data including uni-, bi-, and multi-directional flow [29, 30, 31].
In this work, we present an improved version of the pedestrian contact interaction trajectory analysis studied in Ref. [28]. While Ref. [28] uses one test statistic for contact interaction trajectory characterization, this study applies a set of test statistics in an attempt to systematically characterize contact interaction trajectories for different pedestrian flow conditions.
The remainder of this paper is organized as follows. Section 2 presents our approach for contact interaction trajectory classification in line with statistical testing approach. In that section, we describe the test procedure including synthetic trajectory generation, evaluation, and classification criteria identification. In Section 3, we briefly summarize the dataset used for our analysis and then discuss the findings of our analysis in terms of contact interaction trajectory classification and contact durations.
2 Methods
Our approach for contact interaction trajectory classification consists of two main parts: a statistical testing procedure (steps 1–3) and experimental dataset analysis (steps 4–6). For the statistical testing procedure, we first generate synthetic trajectories by simulating confined diffusion based on Ornstein-Uhlenbeck (OU) process and apply a correlated random walk (CRW) model for ballistic motion (step 1). In step 2, we then evaluate trajectory properties to reflect turning angle movement and the linearity of contact interaction trajectories. In step 3, we quantify classification criteria for the statistical testing procedure. For the experimental dataset analysis, we apply the statistical testing procedure to real data collected from pedestrian flow experiments. In doing so, we import contact trajectories from experimental datasets in step 4, and in step 5 we evaluate the trajectory properties as in step 2. In step 6, we then classify the trajectories based on the classification criteria that developed in step 3. A schematic representation of contact interaction trajectory classification can be seen from Fig. 1.
Figure 1: Schematic representation of our approach for contact interaction trajectory classification: (step 1) generate synthetic trajectories based on Ornstein-Uhlenbeck (OU) process for confined motion and applying correlated random walk (CRW) model for ballistic motion, (step 2) evaluate trajectory properties, (step 3) quantify classification criteria for the statistical testing procedure, (step 4) import contact trajectories from experimental datasets, (step 5) evaluate trajectory properties like in step 2, and (step 6) classify trajectories based on the criteria quantified in step 3.
2.1 Generation of synthetic trajectories
We generate a set of synthetic contact interaction trajectories in line with random walk models. Two types of motions are considered for the trajectory generation: confined diffusion and ballistic motion. If a trajectory shows confined diffusion, the trajectory remains in a restricted area [35, 36, 37]. In the context of contact interactions, confined diffusion results in long or even infinite duration of contact interaction among individuals. On the other hand, if a trajectory shows ballistic motion, individuals keep their walking direction in the subsequent steps. Consequently, individuals tend to move in straight lines between their start and end points, likely showing short duration of contact interactions. We apply a stochastic differential equation (given in Eq. (1)) to generate the confined motion trajectories and correlated random walk model for the ballistic motion trajectories.
Table 1: Parameter setup for synthetic trajectory generation.
Motion type Model Parameter
Confined motion OU $n=[10,1000]$
$\lambda=(c,1]$ with $c=0.1$
$\mu=0$
$\sigma=1$
Ballistic motion CRW $n=[10,1000]$
$\kappa=[0,500]$
$\beta_{0}=0$, $\beta(0)=0$
$v=1$
$\Delta t=1$
2.1.1 Ornstein-Uhlenbeck (OU) process
According to Refs. [21, 25, 26], we numerically simulate confined motion based on Ornstein-Uhlenbeck (OU) process, which is given in terms of the following stochastic differential equation:
$dX_{t}=-\lambda(X_{t}-\mu)dt+\sigma dB^{1/2}_{t},$ (1)
where $X_{t}=(x(t),y(t))$ is the position in the two-dimensional space obtained at time $t$. The equilibrium position is given as $\mu=(\mu_{x},\mu_{y})$. The interaction strength is indicated as $\lambda=(\lambda_{x},\lambda_{y})$, and its value follows a uniform distribution. Based on the parameter values presented in Table 1, we simulate 10000 trajectories of the Ornstein-Uhlenbeck (OU) process in Eq. (1). The simulation time step is set to 1 and the speed is controlled by means of the interaction strength $\lambda$. The simulation parameter values are selected in line with previous studies [21, 25, 26].
2.1.2 Correlated random walk (CRW) model
We simulate ballistic motion based on the correlated random walk (CRW) model. For each simulation time step, we update the position of contact interaction trajectory for a given speed $v$:
$\begin{split}x(t+\Delta t)&=x(t)+vcos(\beta(t))\Delta t,\\ y(t+\Delta t)&=y(t)+vsin(\beta(t))\Delta t.\\ \end{split}$ (2)
Here, $\Delta t$ is the time step size and $\beta(t)$ is the turning angle which is given as a
$\beta(t)=\beta(t-\Delta t)+\Delta\beta.$ (3)
In accordance with previous studies [12, 38, 39, 40], the turning angle increment $\Delta\beta$ is sampled from the von Mises distribution $\nu(\beta_{0}|\kappa)$, where $\beta_{0}$ is the mean angle and $\kappa$ is the scale parameter reflecting directional correlation. Note that ballistic motion trajectories can be generated when the value of $\kappa$ is large. In contrast, if $\kappa$ is small, the generated trajectory tends to show pure random walk behavior. For each $\kappa$, we generate synthetic trajectories of ballistic motions by performing 10000 independent simulation runs. The time step $\Delta t$, is set as 1 and the speed $v$ is 1, so the walker moves a unit length at each time step. The simulation parameter values in Table 1 are selected based on Refs. [12, 39].
2.2 Trajectory characterization
We characterize contact interaction trajectories based on two test statistics: turning angle entropy $H$ and efficiency $E$.
2.2.1 Turning angle entropy
Figure 2: Schematic representation of turning angle. We compute the turning angle $\theta(t)$ at $B$ based on three successive positions $A$, $B$, and $C$, which represent the positions $[x(t-\tau),y(t-\tau)]$, $[x(t),y(t)]$, and $[x(t+\tau),y(t+\tau)]$, respectively. Here, $\varphi_{AB}$ is the angle of vector $\overrightarrow{AB}$ relative to positive $x$ axis, and $\varphi_{BC}$ for $\overrightarrow{BC}$. Note that $\theta$ is considered positive in counterclockwise direction.
To understand the change of movement direction in contact interaction trajectories, we firstly obtained the turning angle $\theta(t)$. Similar to previous studies [41, 42, 43], we compute the turning angle $\theta(t)$ based on three successive positions $[x(t-\tau),y(t-\tau)]$, $[x(t),y(t)]$, and $[x(t+\tau),y(t+\tau)]$, which are denoted by $A$, $B$, and $C$ in Fig. 2. The sampling interval $\tau$ determines how often we sample data points from a contact interaction trajectory to evaluate $\theta(t)$.
The turning angle $\theta(t)$ indicates the change in walking direction from the previous movement $\overrightarrow{AB}$ to the current movement $\overrightarrow{BC}$, i.e., $\varphi_{AB}$ and $\varphi_{BC}$ respectively. We calculate the angle $\varphi_{AB}$ of vector $\overrightarrow{AB}$ relative to positive $x$ axis:
$\varphi_{AB}=k_{AB}\pi+arctan\left[\frac{y(t)-y(t-\tau)}{x(t)-x(t-\tau)}\right],$ (4)
with
$k_{AB}=\begin{cases}0,&\text{if $x(t)>x(t-\tau)$ and $y(t)>y(t-\tau)$}\\ 1,&\text{if $x(t)<x(t-\tau)$}\\ 2,&\text{if $x(t)>x(t-\tau)$ and $y(t)<y(t-\tau)$}.\\ \end{cases}$
Likewise, we calculate the angle $\varphi_{BC}$ of vector $\overrightarrow{BC}$ relative to positive $x$ axis:
$\varphi_{BC}=k_{BC}\pi+arctan\left[\frac{y(t+\tau)-y(t)}{x(t+\tau)-x(t)}\right],$ (5)
with
$k_{BC}=\begin{cases}0,&\text{if $x(t+\tau)>x(t)$ and $y(t+\tau)>y(t-\tau)$}\\ 1,&\text{if $x(t+\tau)<x(t)$}\\ 2,&\text{if $x(t+\tau)>x(t)$ and $y(t+\tau)<y(t)$}.\\ \end{cases}$
Based on Equations 4 and 5, the turning angle $\theta(t)$ is given as
$\theta(t)=2m\pi+\varphi_{BC}-\varphi_{AB},$ (6)
with
$m=\begin{cases}0,&\text{if $\left|\varphi_{BC}-\varphi_{AB}\right|<\pi$}\\ -1,&\text{if $\left|\varphi_{BC}-\varphi_{AB}\right|>\pi$ and $\varphi_{BC}>\varphi_{AB}$}\\ 1,&\text{if $\left|\varphi_{BC}-\varphi_{AB}\right|<\pi$ and $\varphi_{BC}<\varphi_{AB}$}.\\ \end{cases}$
Note that the range of turning angle $\theta(t)$ is between $-\pi$ and $\pi$.
For each contact interaction, we then define Shannon entropy of turning angle $H$ as
$H=-\sum_{i}^{N}p_{i}(\theta)\log_{N}p_{i}(\theta),$ (7)
where $N$ is the number of divided bins in the complete range of turning angle and $p_{i}(\theta)$ is the probability that a turning angle $\theta$ is observed in bin $i$. In this study, the complete range of turning angles $\theta\in[-\pi,\pi]$ are evenly divided into $N=24$ sections with a constant size of $\Delta\theta=\pi/12=15^{\circ}$. We compute the probability $p_{i}(\theta)$ by counting the frequency of $\theta$ within each bin. The turning angle entropy $H$ lies between 0 and 1. When the contact interaction trajectory is a straight line, the turning angle entropy $H$ is 0, indicating that the turning angle is not changing during the contact interaction. In case of $H=1$, the distribution of turning angle is uniform, indicating that every angle (bin) is observed, thus we can see that the turning angle is changing frequently.
2.2.2 Efficiency
Efficiency $E$ measures the ratio of the net squared displacement of a trajectory to the sum of squared step lengths, reflecting the linearity of the trajectory. For a trajectory containing $n$ data points of position, Efficiency $E$ is defined as
$E=\frac{|X_{n-1}-X_{0}|^{2}}{\sum_{i=1}^{n-1}|X_{i}-X_{i-1}|^{2}},$ (8)
where $X_{i}$ denotes the position at time instance $i$ and $X_{0}$ for the position at the start of contact interaction.
2.3 Classification criteria
In line with the statistical test procedure presented in previous studies [21, 22, 25], we identify classification criteria for different motion types based on the knowledge of turning angle entropy $H$ and efficiency $E$ distributions that we can observe from the synthetic trajectories.
Figure 3: The average end-to-end distance $\left\langle L\right\rangle$ as a function of average trajectory length $\left\langle l\right\rangle$ measured from: (a) correlated random walk (CRW) trajectories and (b) Ornstein-Uhlenbeck (OU) process trajectories.
First, we evaluate the scaling behavior of average end-to-end distance $\left\langle L\right\rangle$ as a function of average trajectory length $\left\langle l\right\rangle$. By doing that, we can quantify the confined and ballistic motions from OU process and CRW model simulations, respectively. Previous studies [45, 46] suggested that one can classify a random walk motion as a ballistic motion if the slope of $\left\langle L\right\rangle$ the curve in a log–log plot is approximately equal to 1, i.e., $\left\langle L\right\rangle\sim\left\langle l\right\rangle$. In addition, it is also suggested that $\left\langle L\right\rangle$ the curve reaches a plateau in the case of confined motion. That is, the trajectory of contact interaction is trapped in a small area, thus the end-to-end distance $L$ does not increase further, even though the trajectory length $l$ increases. Figure 3 shows a log–log plot of average end-to-end distance $L$ as a function of average trajectory length $l$ for the CRW model and OU process simulations. Analogous to previous studies [22, 25], we estimate the slope of $\left\langle L\right\rangle$ curve by fitting a function $a\log(t)+b$ to the estimated $\log(\left\langle L\right\rangle)$ curve. We choose the cutoff value of the exponent as $c=0.05$. When the slope of $\left\langle L\right\rangle$ curve in a log–log plot is between $1.951-c=$ 1.95 and $2.051+c=$ 2.05, i.e., $a\in[1.95,2.05]$, we consider the corresponding trajectory shows ballistic motion. If the value of $a$ is near 0, i.e., $a\in[-0.05,0.05]$, the corresponding trajectory is classified as confined motion. In line with the scaling behavior of $\left\langle L\right\rangle$ presented in Fig. 3, we categorize synthetic trajectories as the ballistic motion if the trajectories are generated by Eq. (2) with $\kappa\geq 300$, and trajectories generated by Eq. (1) as the confined motion.
Figure 4: A representative example of turning angle entropy $H$ and efficiency $E$ distributions for the ballistic motion: (a) relative frequency and (b) cumulative distribution. A sample trajectory is shown in (c). The red shaded area in (b) indicates the critical region with the significance level $\alpha=0.05$ suggested in the previous studies Refs. [20, 21, 25]. Distribution curves are generated based on 10000 independent simulation runs with $\kappa=300$ and $n=300$.
Figure 5: Same as Fig. 4 but for the confined motion. Distribution curves are generated based on 10000 independent simulation runs with $n=300$.
Next, we define a critical region of turning angle entropy $H$ and efficiency $E$ distributions for the ballistic and confined motions. For different numbers of (positional) data points, we estimate the boundaries of the critical region for test statistics $H$ and $E$. For the simulated trajectories generated by the CRW model with $\kappa\geq 300$, we assume that ballistic motion will tend to show low $H$ and high $E$. Figure 4 shows a representative example of turning angle entropy $H$ and efficiency $E$ distributions for ballistic motion. As can be seen from Table 2, the upper boundary of $H$ is identified as
$H\leq q_{1-\alpha},$ (9)
where $q_{1-\alpha}$ indicates that $H$ lies in the critical region with the probability $1-\alpha$. The lower boundary of $E$ is identified as
$q_{\alpha}\leq E,$ (10)
where $q_{\alpha}$ indicates that $E$ lies in the critical region with the probability $1-\alpha$. We use $\alpha=0.05$ according to Refs. [21, 25]. One can notice that the turning angle entropy $H$ is stable, while the efficiency $E$ is notably decreasing as the number of data points $n$ increases. During the ballistic motion, individuals tend maintain their walking direction in the subsequent steps, thus the change in walking direction between two consecutive steps is small, yielding low $H$. As the individuals continue walking (i.e., $n$ is increasing), sometimes the change in walking direction adds up, resulting in deviation from a straight line and consequently leading to decreasing $E$ for increasing $n$. Thus, it is suggested that the turning angle entropy $H$ is a useful measure for the ballistic motion.
In contrast to the ballistic motion trajectories, the confined motion trajectories generated using the OU process yield high $H$ and low $E$ as shown in Fig. 5. Similarly to Eqs. (9) and (10), we identify the lower boundary of $H$ as
$q_{\alpha}\leq H,$ (11)
and the upper boundary of $E$ as
$E\leq q_{1-\alpha}.$ (12)
In Table 3, the turning angle entropy $H$ is increasing rapidly for the growing number of data points $n$ particularly for $n<100$, but the efficiency $E$ is virtually constant for a wide range of $n$. During the confined motion, individuals continue changing walking direction and the number of data points $n$ is increasing. When $n$ is small, the turning angle entropy $H$ might be underestimated due to the limited number of data points. In this case, the value of $H$ is increasing as $n$ grows, which can be understood that $H$ is reaching the stationary state value once $n$ becomes sufficiently large. Meanwhile, $E$ is stable for a wide range of $n$, suggesting that the efficiency $E$ is a good measure for the confined motion.
Table 2: Classification criteria for ballistic motion. For different number of data points $n$, we evaluate the critical region in the distributions of turning angle entropy $H$ and efficiency $E$ with the significance level $\alpha=0.05$.
number of data points $n$
test statistics 10 25 50 100 500 1000
entropy $H$ smaller than 0.21 0.22 0.22 0.22 0.26 0.26
efficiency $E$ larger than 0.98 0.96 0.91 0.82 0.33 0.11
Table 3: Same as Table 2 but for the confined motion.
number of data points $n$
test statistics 10 25 50 100 500 1000
entropy $H$ larger than 0.21 0.53 0.68 0.78 0.89 0.90
efficiency $E$ smaller than 0.09 0.05 0.05 0.05 0.05 0.05
Figure 6: Summary of classification criteria: (a) a classification tree (b) a schematic diagram presenting the criteria in $(H,E)$ space.
To identify classification criteria applicable over a wide range of $n$, we use turning angle entropy $H$ and efficiency $E$ to characterize the ballistic and confined motions, respectively. Consequently, the ballistic motion is characterized by
$H\leq 0.26,$ (13)
and the confined motion is characterized by
$E\leq 0.09.$ (14)
In addition, we define sub-ballistic motion for the remainder of the parameter space, which does not belong to either ballistic or confined motions. Figure 6 summarizes the presented classification criteria.
3 Results and Discussion
3.1 Datasets
Figure 7: Schematic representation of experiment setups: (a) unidirectional flow (scenario name: uni-05), (b) bidirectional flow (scenario name: bi-b01), (c) 2-way crossing flow (scenario name: crossing-90-d08), (d) 3-way crossing flow (scenario name: crossing-120-b01), and (e) 4-way crossing flow (scenario name: crossing-90-a10). Here, blue thick arrows show the walking direction of incoming pedestrians entering the corridors and red thin arrows indicates the walking direction of outgoing pedestrians leaving the corridors. A pair of gray rectangles is placed to set up an entrance of pedestrian group entering the corridor.
Figure 7 shows the sketches of the various experimental setups: uni-directional flow, bi-directional flow, 2-way crossing flow, 3-way crossing flow, and 4-way crossing flow. In the uni-directional flow setup, pedestrians were walking to the right in a straight corridor of 5 m wide and 18 m long. In the bi-directional flow case, two groups of pedestrians enter a straight corridor of 4 m wide and 10 m long through 4 m wide entrance and then walking opposite directions. They leave the corridor through the open passage once they reached the other side of the corridor. In the 2-way, 3-way, and 4-way crossing flows, different groups of pedestrians enter the corridor through 4 m wide entrance and walked 5 m before and after passing through an intersection (4 m by 4 m rectangle in 2-way crossing and 4-way crossing flows, and 4 m wide equilateral triangle in 3-way crossing flow). Similar to the setup of the bi-directional flow, pedestrian groups leave the corridor through the open passage at the end of the corridors. Trajectories of the bi-directional, 2-way, and 3-way crossing flows were recorded at 16 frames per second (fps) and 25 fps for uni-directional and 4-way crossing flow. A more detailed description of the experiment setups can be found in Refs. [29, 30, 31].
From the experimental data, we extracted pairs of individuals in contact and their relative motion trajectories. We consider a pair of individuals to be in contact when the two individuals are within a contact radius. The contact radius $r_{c}$ would depend on the form of transmission in question. In this paper, we assume a 2 m radius based on previous studies [12, 32, 33, 34]. For the analysis, we consider contact interaction trajectories with at least 0.5 seconds records, i.e., 8 data points for the bi-directional, 2-way, and 3-way crossing flows, and 12 data points for the uni-directional and 4-way crossing flow based on literature [21, 25]. Table 4 shows the basic statistics of the selected scenarios, including the number of individuals $N$, experiment period, and the number of contacts. In this study, we consider three types of contact interactions: parallel, head-on, and crossing contact interactions. When two individuals in contact are moving in the same direction, one can define the parallel contact interaction. A head-on contact interaction is observed when two individuals are moving toward opposite directions. We can see a crossing contact interaction from a pair of individuals in contact when one individual is moving perpendicular to the other’s moving direction. The number of contacts is given as the number of interacting individual pairs. In Appendix B, we present the basic descriptive statistics of the other experimental scenarios.
Table 4: Basic descriptive statistics of representative scenarios.
No. contacts
Setup fps Scenario name N Period (s) Parallel Head-on Crossing total
Uni-directional 25 uni-05 905 157.68 25260 (100%) 0 0 25260
Bi-directional 16 bi-b10 736 324.56 17154 (29.81%) 40398 (70.19%) 0 57552
2-way crossing 16 crossing-90-d08 592 147.88 22418 (53.39%) 0 19574 (46.61%) 41992
3-way crossing 16 crossing-120-b01 769 215.63 34654 (36.80%) 59518 (63.20%) 0 94172
4-way crossing 25 crossing-90-a10 324 94.08 882 (9.60%) 3834 (41.73%) 4472 (48.67%) 9188
3.2 Classification results
Figure 8: Representative types of pedestrian contact interaction trajectories characterized based on the value of turning angle entropy $H$ and efficiency $E$ (refer to Eqs. (13) and (14), and Fig. 6): (a) ballistic motion with $H=0.21$ and $E=0.99$ (individual id 65 and id 99 in 4-way crossing scenario crossing-90-a10), (b) confined motion with $H=0.87$ and $E=0.02$ (individual id 25 and id 46 in unidirectional flow scenario uni-05), and (c) sub-ballistic motion with $H=0.71$ and $E=0.38$ (individual id 463 and id 566 in bi-directional scenario bi-b10). The position of focal individuals (id 25 in (a), id 65 in (b), and id 463 in (c)) is indicated at (0, 0) by a black cross symbol $\times$. Red dashed circles show contact range of a focal individual ($r_{c}=2$ m). The relative motion of pedestrians interacting with the focal individuals (id 46 in (a), id 99 in (b), and id 566 in (c)) denoted by blue solid lines. In the lower panels, blue dashed lines indicate the ground truth trajectories of focal individuals (id 65 shown in (a), id 25 in (d), and id 463 in (c)), and red solid lines for pedestrians interacting with the focal individuals (id 99 shown in (a), id 46 in (b), and id 566 in (c)). Arrows are guide for the eyes, indicating the walking direction of individuals.
Figure 8 presents representative types of pedestrian contact interaction trajectories characterized by turning angle entropy $H$ and efficiency $E$ (see Eqs. (13) and (14), and Fig. 6)). As can be seen from Fig. 8(a), the relative motion of individual $j$ moves in parallel along a straight line, showing a ballistic motion. In Fig. 8(b), the relative motion trajectory of individual $j$ stays close to its initial position during the contact interaction, implying confined motion. It should be noted that confined motion in contact interaction trajectories does not necessarily suggest that ground truth trajectories display confined motions. In the case of sub-ballistic motion (see Fig. 8(c)), the relative motion of individual $j$ occasionally shows changes in its walking direction but still travels far away from its initial position during the contact interaction.
Figure 9: Summary of motion type classification results. For each panel, horizontal rows correspond to contact interaction types and vertical columns to the motion types.
Based on the classification criteria presented in section 2.3, we have classified motion types of contact interaction trajectories. The summary of the motion type classification results for the selected scenarios can be found in Fig. 9. We can observe that ballistic motion is more frequently observed in the presented scenarios of the bi-directional (scenario bi-b10) and 4-way crossing flow (scenario crossing-90-a10). In contrast, the majority of contact interactions in the unidirectional flow setup (scenario uni-05) are categorized as confined motion, hinting at the possibility of long-lived contact duration. This implies that the risk level of virus exposure in human face-to-face intersections is high. In addition, for other experiment scenarios including crossing-90-d08 (2-way crossing setup) and crossing-120-b01 (3-way crossing setup), the sub-ballistic motion is the most frequently observed motion type. A summary of motion type classification results for other experiment scenarios are given in Appendix C.
Figure 10: Average values of contact duration $t_{c}$ measured for different contact trajectory type classification. For each panel, horizontal rows correspond to contact interaction types and vertical columns to the motion types. Dark green cell indicates high value of $t_{c}$.
Figure 10 illustrates average values of contact duration $t_{c}$ measured for different contact trajectory type classification. For the presented scenarios, confined motions observed from parallel contact interaction tend to yield high value $t_{c}$, displaying striking difference from $t_{c}$ measured from ballistic motions. This tendency can be also seen from other scenarios, refer to Appendix D.
Figure 11: Relationship among the parallel contact interactions, confined motion, and contact duration $t_{c}$: (a) the frequency of confined motion increases as that of parallel contact interaction grows, (b) the average value of contact duration $t_{c}$ measured for the confined motions tends to increase as more confined motions are observed from the experiments, and (c) correlation matrix showing. In (a) and (b), different symbols indicate different experiment setup. In (c), "tc_all" indicates the average contact duration evaluated for each experiment scenario regardless of motion types. The frequency of parallel contact interactions and confined motions are indicated by "parallel" and "confined", respectively.
To understand the influence of confined motions observed from parallel contact interaction on the contact duration $t_{c}$, we take a closer look at the relationship among the parallel contact interactions, confined motion, and $t_{c}$, see Fig. 11. As in shown Fig. 11(a), the frequency of confined motion increases as the frequency of parallel contact interaction grows. One can see from Fig. 11(b) that the average value of contact duration $t_{c}$ measured for the confined motions shows an increasing trend when the confined motions are more frequently observed from the experiments. Based on Fig. 11(a) and Fig. 11(b), it is suggested that more parallel contact interactions tend to produce more confined motions, and consequently yields longer contact duration $t_{c}$. That is compatible with high correlation among the parallel contact interactions, confined motion, and contact duration $t_{c}$ presented in Fig. 11(c).
4 Conclusion
In this study, we have examined how the pedestrian flow setups affect contact duration distribution which is highly relevant to the spreading risk in human crowds. To this end, we have analyzed pedestrian contact interaction trajectories of uni-, bi- and multi-directional flow setups in experimental data [29, 30, 31]. In line with statistical testing approach, we have classified types of motions observed in the contact interactions based on the turning angle entropy $H$ and efficiency $E$. We generated synthetic trajectories and then evaluated the trajectory properties to systematically quantify classification criteria. The identified criteria were then applied to classify the contact interaction trajectories collected from pedestrian flow experiments.
From the experimental dataset, we have classified the pedestrian contact interaction trajectories for three types of motions: ballistic, confined, and sub-ballistic motions. We observed the ballistic motion when $H$ is low. In the case of ballistic motion, the contact interaction trajectories move in parallel along a straight line, so the contact duration tends to be brief. On the other hand, the confined motion can be defined when $H$ is high and $E$ is low. This indicates that the contact interaction trajectories stay close to their initial position during the contact interaction. From the experimental data of various pedestrian flow setups, we have found that confined motion appears more as the frequency of parallel contact interactions increases. It is noted that the confined motions tend to yield longer contact duration $t_{c}$, potentially leading to higher risk of virus exposure in human face-to-face interactions.
A few selected experimental setups have been analyzed to study the fundamental role of pedestrian flow setups (e.g., uni-, bi-, and multi-directional flow) in the distribution of pedestrian motion types and contact duration. To generalize the findings of this study, the presented analysis should be further performed with larger number of scenarios and various layouts of pedestrian facilities at different locations. Another interesting extension of the presented study can be planned in line with the importance analysis of factors influencing on the motion types and duration of contact interactions [22, 27, 49, 50, 51].
Acknowledgements
This research is supported by Ministry of Education (MOE) Singapore under its Academic Research Fund Tier 1 Program Grant No. RG12/21 MoE Tier 1.
Appendix A Data and code availability
The data used for our study can be found from https://ped.fz-juelich.de/da/doku.php?id=start#data_section (last accessed 3 January, 2024). Table 5 shows short URL addresses where readers can find the descriptions and trajectory files of the presented experiment setups.
Table 5: Short URL addresses of the presented experiment setups (last accessed 3 January, 2024).
Setup Short URL
Uni-directional http://ped.fz-juelich.de/da/2013unidirectional
Bi-directional http://ped.fz-juelich.de/da/2013bidirectional
2-way crossing http://ped.fz-juelich.de/da/2013crossing90
3-way crossing http://ped.fz-juelich.de/da/2013crossing120
4-way crossing http://ped.fz-juelich.de/da/2013crossing90
The data processing, simulations, and analysis were carried out in Python with the help of open-source libraries. The code used for our study is publicly available at https://doi.org/10.5281/zenodo.10455825.
Appendix B Basic descriptive statistics of experiment scenarios
In this section, we show basic descriptive statistics of experiment scenarios same as Table 4 for different pedestrian flow setups, see Table 6.
Table 6: Basic descriptive statistics of experiment scenarios.
No. contacts
Setup fps Scenario name N Period (s) Parallel Head-on Crossing total
Uni-directional 25 uni-01 148 75.52 644 (100%) 0 0 644
Uni-directional 25 uni-02 760 191.68 8220 (100%) 0 0 8220
Uni-directional 25 uni-03 916 161.84 17396 (100%) 0 0 17396
Uni-directional 25 uni-04 909 163.28 23214 (100%) 0 0 23214
Uni-directional 25 uni-05 905 157.68 25260 (100%) 0 0 25260
Uni-directional 25 uni-06 913 170.76 27948 (100%) 0 0 27948
Uni-directional 25 uni-07 914 205.08 34696 (100%) 0 0 34696
Uni-directional 25 uni-08 477 118.04 22164 (100%) 0 0 22164
Uni-directional 25 uni-09 310 78.92 16378 (100%) 0 0 16378
Bi-directional 16 bi-a1 377 95.38 5438 (26.86%) 14810 (73.14%) 0 20248
Bi-directional 16 bi-a2 522 143.94 13460 (30.20%) 31114 (69.80%) 0 44574
Bi-directional 16 bi-a3 542 143.00 15718 (30.22%) 36302 (69.78%) 0 52020
Bi-directional 16 bi-a4 544 165.31 18086 (30.09%) 42024 (69.91%) 0 60110
Bi-directional 16 bi-a5 560 177.31 17190 (26.63%) 47372 (73.37%) 0 64562
Bi-directional 16 bi-b01 141 118.88 386 (21.40%) 1418 (78.60%) 0 1804
Bi-directional 16 bi-b02 259 145.56 1148 (20.29%) 4510 (79.71%) 0 5658
Bi-directional 16 bi-b03 480 202.88 4088 (24.33%) 12712 (75.67%) 0 16800
Bi-directional 16 bi-b04 743 296.50 15586 (25.04%) 46654 (74.96%) 0 62240
Bi-directional 16 bi-b05 643 269.75 14192 (22.43%) 49080 (77.57%) 0 63272
Bi-directional 16 bi-b06 830 388.81 24804 (26.02%) 70524 (73.98%) 0 95328
Bi-directional 16 bi-b07 606 254.56 19208 (25.71%) 55508 (74.29%) 0 74716
Bi-directional 16 bi-b08 703 359.69 23202 (24.44%) 71738 (75.56%) 0 94940
Bi-directional 16 bi-b09 483 186.38 6426 (28.54%) 16088 (71.46%) 0 22514
Bi-directional 16 bi-b10 736 324.56 17154 (29.81%) 40398 (70.19%) 0 57552
2-way crossing 16 crossing-90-d01 603 199.06 5248 (27.74%) 0 13672 (72.26%) 18920
2-way crossing 16 crossing-90-d02 604 192.00 6722 (46.34%) 0 7784 (53.66%) 14506
2-way crossing 16 crossing-90-d03 606 186.69 2674 (24.79%) 0 8114 (75.21%) 10788
2-way crossing 16 crossing-90-d05 600 153.75 5404 (26.06%) 0 15330 (73.94%) 20734
2-way crossing 16 crossing-90-d06 597 131.56 15582 (51.03%) 0 14952 (48.97%) 30534
2-way crossing 16 crossing-90-d07 604 139.50 10568 (28.48%) 0 26544 (71.52%) 37112
2-way crossing 16 crossing-90-d08 592 147.88 22418 (53.39%) 0 19574 (46.61%) 41992
3-way crossing 16 crossing-120-a1 254 78.88 1156 (21.85%) 4134 (78.15%) 0 5290
3-way crossing 16 crossing-120-a2 286 93.56 1428 (27.20%) 3822 (72.80%) 0 5250
3-way crossing 16 crossing-120-a3 341 96.56 3638 (26.57%) 10054 (73.43%) 0 13692
3-way crossing 16 crossing-120-a4 710 148.69 14714 (26.64%) 40514 (73.36%) 0 55228
3-way crossing 16 crossing-120-a5 814 186.12 27026 (27.41%) 71578 (72.59%) 0 98604
3-way crossing 16 crossing-120-a6 783 182.44 29764 (30.62%) 67456 (69.38%) 0 97220
3-way crossing 16 crossing-120-a7 886 205.06 33904 (28.90%) 83406 (71.10%) 0 117310
3-way crossing 16 crossing-120-b1 769 215.63 34654 (36.80%) 59518 (63.20%) 0 94172
3-way crossing 16 crossing-120-b2 700 182.38 29636 (37.93%) 48492 (62.07%) 0 78128
3-way crossing 16 crossing-120-c1 262 86.81 1758 (36.76%) 3024 (63.24%) 0 4782
3-way crossing 16 crossing-120-c2 382 99.12 7184 (36.38%) 12564 (63.62%) 0 19748
3-way crossing 16 crossing-120-c3 573 156.75 18016 (39.00%) 28176 (61.00%) 0 46192
3-way crossing 16 crossing-120-c4 697 196.75 28546 (39.33%) 44040 (60.67%) 0 72586
3-way crossing 16 crossing-120-c5 627 201.81 30788 (40.41%) 45402 (59.59%) 0 76190
3-way crossing 16 crossing-120-c6 811 274.50 38872 (39.89%) 58574 (60.11%) 0 97446
3-way crossing 16 crossing-120-c7 842 191.50 54378 (42.60%) 73270 (57.40%) 0 127648
4-way crossing 25 crossing-90-a01 247 168.08 68 (3.85%) 920 (52.04%) 780 (44.12%) 1768
4-way crossing 25 crossing-90-a02 439 132.28 3186 (15.66%) 7566 (37.19%) 9594 (47.15%) 20346
4-way crossing 25 crossing-90-a03 323 90.76 2816 (15.18%) 6718 (36.22%) 9014 (48.60%) 18548
4-way crossing 25 crossing-90-a04 352 105.04 4636 (13.61%) 11712 (34.38%) 17716 (52.01%) 34064
4-way crossing 25 crossing-90-a05 337 88.08 6210 (16.83%) 11912 (32.29%) 18770 (50.88%) 36892
4-way crossing 25 crossing-90-a06 269 47.56 3234 (13.12%) 6150 (24.96%) 15260 (61.92%) 24644
4-way crossing 25 crossing-90-a07 323 100.04 2338 (15.60%) 5564 (37.13%) 7082 (47.26%) 14984
4-way crossing 25 crossing-90-a08 298 71.96 7850 (25.11%) 7482 (23.93%) 15928 (50.95%) 31260
4-way crossing 25 crossing-90-a09 299 116.24 310 (7.03%) 2066 (46.83%) 2036 (46.15%) 4412
4-way crossing 25 crossing-90-a10 324 94.08 882 (9.60%) 3834 (41.73%) 4472 (48.67%) 9188
Appendix C Summary of motion type classification results
In this section, we present summary of motion type classification results same as Fig. 9 for different pedestrian flow setups: uni-directional (Fig. 12), bi-directional (Fig. 13), 2-way crossing (Fig. 14), 3-way crossing (Fig. 15), and 4-way crossing flows (Fig. 16).
Figure 12: Summary of motion type classification results for unidirectional flow scenarios.
Figure 13: Summary of motion type classification results for bidirectional flow scenarios.
Figure 14: Summary of motion type classification results for 2-way crossing flow scenarios.
Figure 15: Summary of motion type classification results for 3-way crossing flow scenarios.
Figure 16: Summary of motion type classification results for 4-way crossing flow scenarios.
Appendix D Average contact duration
In this section, we illustrate average values of contact duration $t_{c}$ same as Fig. 10 for different pedestrian flow setups: uni-directional (Fig. 17), bi-directional (Fig. 18), 2-way crossing (Fig. 19), 3-way crossing (Fig. 20), and 4-way crossing flows (Fig. 21).
Figure 17: Average values of contact duration $t_{c}$ for unidirectional flow scenarios.
Figure 18: Average values of contact duration $t_{c}$ for bidirectional flow scenarios.
Figure 19: Average values of contact duration $t_{c}$ for 2-way crossing flow scenarios.
Figure 20: Average values of contact duration $t_{c}$ for 3-way crossing flow scenarios.
Figure 21: Average values of contact duration $t_{c}$ for 4-way crossing flow scenarios.